# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ 14. The Basics: $KO$-theory for Spaces
. For a compact space $X$, define
@@
    so, in particular, $TS^{n}$ plus a trivial bundle is isomorphic to a trivial bundle, whether $TS^{n}$ is trivial or not. Such a bundle is called stably trivial. So the tangent bundle of $S^{n}$ is stably trivial for all $n$. It is a rather deep result that $TS^{n}$ is isomorphic to a trivial bundle if and only if n=1, 3,n=1,\,3, or $7$. (This is related to the existence of the complex numbers, the quaternions, and the Cayley numbers in dimensions 2, 4,2,\,4, and $8$ and the fact that there are no other such algebras in other dimensions.)
@@
$Vect^{\mathbb{R}}(X)\simeq\it{Proj^{R}}(C^{\mathbb{R}}(X))$
between the category $Vect^{\mathbb{R}}(X)$ of isomorphism classes of vector bundles over $X$ and the category $\it{Proj^{R}}(C^{\mathbb{R}}(X))$ of projections in matrix algebras over $C^{\mathbb{R}}(X)$. This induces a natural isomorphism of functors
@@
called Chern classes (cf. Milnor-Stasheff [63], Chapter 14). These may be defined using differential geometry or purely topologically. Here’s a quick sketch. Suppose $E\to X$ is a complex bundle of dimension $n$. We use the fact that the integral cohomology of the space $BU_{n}$ is a polynomial algebra over the integers with generators $\{c_{1},\dots,c_{n}\}$ with $dim(c_{j})=2j$:
@@
    $ch:K^{*}(X)\otimes{\mathbb{Q}}\overset{\cong}{\longrightarrow}\bigoplus_{k}H^{**}(X;{\mathbb{Q}}).$
@@
Moving on, we come to the case of $KO^{*}(X)$, and for this we refer to Milnor-Stasheff [63] Chapter 15 and to Hatcher [40]. If $E\to X$ is a real vector bundle with complexification $E\otimes{\mathbb{C}}$. It turns out that
@@
    $Ph:KO^{*}(X)\longrightarrow\bigoplus_{k}H^{**}(X;{\mathbb{Q}}).$
@@
(Grady and Sati [38], §2.1.) It’s interesting to look at the simplest case, when $X$ is a point $pt$. Then $KO^{-*}(pt)=KO_{*}({\mathbb{R}})$ and so $KO^{*}(pt)\otimes{\mathbb{Q}}={\mathbb{Q}}$ when =0,4*=0,4 and is 0 otherwise. The generator of $KO^{0}(pt)$ is the identity of the ring, and is represented by a one-dimensional trivial bundle over the point. It goes to the generator of $H^{0}(pt;{\mathbb{Z}})\cong{\mathbb{Z}}\subset H^{0}(pt;{\mathbb{Q}})$ since $Ph$ is a ring map. The group $KO^{-4}(pt)$ is more interesting. We have a natural isomorphism
@@
15. Relating $K$-theory for $C^{*}$-Algebras and $K$-theory for Spaces
@@
The map $\Theta$ is an isomorphism.
@@
defines an injective map of vector bundles from $E$ to $X\times{\mathbb{R}}^{nk}$. Using the standard inner product on ${\mathbb{R}}^{nk}$, for each $x\in X$, let $p_{x}$ be the orthogonal projection from ${\mathbb{R}}^{nk}$ to the range of $f$ over $x$. Then $p_{x}$ is a continuous function of $x$ and is a projection in the algebraic sense that $p^{2}=p$ and $p^{*}=p$. Thus $p]\in\mathcal{P}_{\infty}(A)$ and by construction we have $\Theta([p])\cong E$. ∎
@@
$KO^{0}(X)=\ker\left[KO^{0}(X^{+})\xrightarrow{\iota^{*}}KO^{0}(+)\right]\;$ (1)
@@
$KO_{0}(A)=\ker\left[KO_{0}(A^{+})\xrightarrow{\pi_{*}}KO_{0}({\mathbb{R}})\right]\;$ (2)
@@
specifically evaluation at the point ++. This evaluation map is dual to the inclusion map $\iota$. This shows that the maps shown in equations (1) and (2) are identical and hence there is an isomorphism $KO^{0}(X)\cong KO_{0}(A)$.
@@
Note that in (2) above, $KR^{0}(~)$” refers to Atiyah’s “Real $K$-theory” which is described in the next section.
@@ 16. Atiyah’s $KR$- theory
We assert that $KR^{*}$-theory for spaces with involution is determined by $KO_{*}$- theory for an associated real $C^{*}$-algebra. Here is the precise statement.
@@
One should compare Karoubi [55] Theorem 4.1 for a related result.
Before beginning the proof, we establish a basic property of complex involutive bundles over spaces with involution.
@@
Let $(X,\tau)$ be a compact space with involution and let $(E,\sigma)$ be a complex vector bundle over $X$ with a conjugate linear involution satisfying $\sigma(E_{x})=E_{\tau(x)}$. Then $(E,\sigma)$ is isomorphic (as a complex vector bundle with involution) to a sub-bundle $(E^{\prime},\sigma )$ of the trivial bundle $(X\times{\mathbb{C}}^{n},\tau\times c_{n})$ for some $n$ (where $c_{n}$ is complex conjugation on ${\mathbb{C}}^{n}$). Further, there is a subbundle $(F,\rho)$ of the trivial bundle such that
$(E^{\prime},\sigma^{\prime})\oplus(F,\rho)\cong(X\times{\mathbb{C}}^{n},\tau\times c_{n})$
@@
First we work locally. We claim that for all $x_{0}\in X$ there exists a neighborhood $\mathcal{U}$ of $x_{0}$ in $X$ such that $\tau(\mathcal{U})=\mathcal{U}$ and $(E_{\mathcal{U}},\sigma)$ is isomorphic to $(\mathcal{U}\times{\mathbb{C}}^{n},\tau\times c_{n})$
Let $x_{0}\in X$ be given and let $\mathcal{U}$ be a neighborhood of $x_{0}$ on which $E$ is trivial. If $\tau(x_{0})\neq x_{0}$ we can arrange without loss of generality that $\mathcal{U}$ is a disjoint union of $\mathcal{V}$ and $\tau(\mathcal{V})$. Let $\phi\colon E_{\mathcal{V}}\rightarrow\mathcal{V}\times{\mathbb{C}}^{n}$ be an isomorphism; we use the notation $\phi_{x}\colon E_{x}\rightarrow{\mathbb{C}}^{n}$ for the map on the fibre. Extend $\phi$ to $E_{\mathcal{\tau(\mathcal{V})}}\rightarrow\tau(\mathcal{V})\times{\mathbb{C}}^{n}$ by
$\phi(\tau(x),v)=(\tau(x),c_{n}(\phi_{x}(\sigma_{\tau(x)}(v))))\;.$
Now $\phi\colon E_{\mathcal{U}}\rightarrow\mathcal{U}\times{\mathbb{C}}^{n}$ is defined and satisfies $(\tau\times c_{n})\circ\phi=\phi\circ\sigma$.
@@
Since $\widetilde{\sigma}_{x}$ is conjugate-linear and depends continuously on $x$, we can write
$\widetilde{\sigma}_{x}=c_{n}\circ A_{x}$
where $A_{x}$ is an $n\times n$ matrix depending continuously on $x$. The condition that $\widetilde{\sigma}$ is an involution implies that ${A_{x}}\overline{A_{\tau(x)}}=I_{n}$ for all $x\in\mathcal{U}$. Indeed, $(\widetilde{\sigma})^{2}=\mathrm{id}$ can be rewritten as
$c_{n}A_{\tau(x)}c_{n}A_{x}v=v\qquad for\qquad v\in{\mathbb{C}}^{n}$
@@
The isomorphism $\psi$ that we are looking for is from $\mathcal{U}\times{\mathbb{C}}^{n}$ to itself, so it can be represented by a continuous matrix-valued function $B_{x}$ for $x\in\mathcal{U}$. We need $\psi$ to carry the involution $\widetilde{\sigma}$ to the involution $c_{n}$ on ${\mathbb{C}}^{n}$. In other words, we need the diagram
@@
$(B_{\tau(x)}\circ c_{n}\circ A_{x})(v)=(c_{n}\circ B_{x})(v).$
@@
We know that there is a continuous square root defined on square matrices in a neighborhood of the identity. We replace $\mathcal{U}$ with a smaller neighborhood to insure that square roots are defined on $A_{x}$ for $x\in\mathcal{U}$, and let $B_{x}=(A_{x})^{1/2}$. Then using the formula $A_{\tau(x)}=(\overline{A_{x}})^{-1}$ we have $B_{\tau(x)}=(\overline{A_{x}})^{-1/2}$ so
@@
$(E,\sigma).\longrightarrow(X\times{\mathbb{C}}^{m},\tau\times c_{n})$
@@
Finally, take $(E,\sigma)=(E,\tau\times c^{n})$ to be a subbundle of the trivial bundle. Then take $F$ to be the point-wise orthogonal complement to $E$ again with the involution $\tau\times c^{n}$. Since vector conjugation on ${\mathbb{C}}^{n}$ respects the inner product we have the statement that
$(\tau\times c^{n})(E)=E\Longrightarrow(\tau\times c^{n})(F)=F.$
Clearly
$(E,\tau\times c^{n})\oplus(F,\tau\times c^{n})=(X\times{\mathbb{C}}^{n},\tau\times c^{n}).$
The rest of this section is devoted to a new proof of Theorem 16.1. Let $(X,\tau)$ be a compact space with involution and let $A=A(X,\tau)$. We shall show that there is a natural isomorphism of abelian semigroups
@@
Proof.
Our proof follows the standard route.
  1. (1)
    Define the map $\Theta$.
  2. (2)
    Show that $\Theta$ is well-defined.
  3. (3)
    Show that $\Theta$ is 1-1.
  4. (4)
    Show that $\Theta$ is onto.
  5. (5)
    Check that $\Theta$ is natural with respect to maps of spaces with involution.
First let $p\in M_{n}(A)$ be a projection. Then $p\in M_{n}(C(X,{\mathbb{C}}))$ and $p(\tau(x))=\overline{p(x)}$. Let
@@
with the associated projection map $\pi\colon E\rightarrow X$ by $\pi(x,v)=x$. Now $E$ is a complex vector bundle and has a conjugate-linear involution $\sigma=\tau\times c_{n}$ given by $(x,v)\mapsto(\tau(x),\overline{v})$. This involution $\sigma=\tau\times c_{n}$ certainly commutes with the respective involutions on $X$ as required. To show that $(\tau(x),\overline{v})\in E$ we check that
@@
We must show that $\Theta$ is well-defined. Suppose that $p$ and $q$ are projections in $M_{n}(A)$ and
$p]=[q]\in\mathcal{P}_{\infty}(A).$
There is a unitary $u\in M_{n}(A)$ such that $upu^{*}=q$. Let
@@
Thus the construction $p\mapsto(E,\sigma)$ defines a homomorphism of abelian semigroups
@@
Next, we claim that $\Theta$ is 111-1. Assume that $p,q$ in $M_{n}(A)$ are projections that satisfy $E_{p}\cong E_{q}$. Similar to the proof of Lemma 15.4, there exists $a\in C(X,M_{n}({\mathbb{C}}))$ where $a_{x}$ is an isomorphism from $p(x){\mathbb{C}}^{n}$ to $q(x){\mathbb{C}}^{n}$ for each $x$. The hypothesis that this isomorphism commutes with the involutions on $E_{p}$ and $E_{q}$ is expressed by the formula $a_{\tau(x)}(\overline{v})=\overline{a_{x}(v)}$, which in turn implies that $a_{\tau(x)}(\overline{v})=\overline{a_{x}}\,\overline{v}$. Hence $a_{\tau(x)}=\overline{a_{x}}$ and so $a\in M_{n}(A)$.
As before, we let $s=f(a^{*}a)a^{*}$ where
@@
noting in particular that because $a\in M_{n}(A)$, we also have $s\in M_{n}(A)$. As before, $s$ is a partial isometry that satisfies $ss^{*}=p$ and $s^{*}s=q$.
Since
@@
we have again that $\mathrm{diag}(p,0)$ is unitarily equivalent to $\mathrm{diag}(q,0)$ in $M_{2n}(A)$, so $[p]=[q]$ in $\rm{Proj}(A)$. Thus $\Theta$ is 1-1.
Next, we show that $\Theta$ is onto. Let $(E,\sigma)$ be any complex vector bundle with involution over $(X,\tau)$. Using Lemma 16.2, identify $E$ with a sub-bundle of $X\times{\mathbb{C}}^{n}$ in such a way that the involution $\sigma$ on $E$ corresponds to conjugation on ${\mathbb{C}}^{n}$. In other words, $\sigma$ can be expressed by the formula
@@
where $\overline{v}$ is component-wise conjugation on ${\mathbb{C}}^{n}$. Define $p(x)\in C(X,M_{n}({\mathbb{C}}))$ by letting $p(x)$ be the orthogonal projection of ${\mathbb{C}}^{n}$ to $E_{x}$. Then $p$ is continuous in $x$, and is a projection: $p^{2}=p^{*}=p$. Furthermore, we claim that our condition on $\sigma$ implies that
$p(\tau(x))=\overline{p(x)}\;.$
Indeed, suppose that $v\in E_{x}$ is in the range of $p(x)$. Then
$\overline{p(\tau(x))}\cdot v=p(x)\cdot v=v\;$
which shows that $v$ is in the range of $\overline{p(\tau(x))}$. As $\overline{p(\tau(x))}$ and p(x){p(x)} are both orthogonal projections with the same rank, they must be equal, implying that $\overline{p(\tau(x))}={p(x)}$ and hence $p(\tau(x))=\overline{p(x)}\;$. From this it follows that $p$ is a projection in $M_{n}(A)$. By construction we have $(E,\sigma)=(E_{p},\overline{~\cdot~})$, which shows that $\Theta$ is surjective.
@@
The proof in the locally compact case follows the same strategy as that in the proof of Theorem 15.2. Note that if $X$ is locally compact then any involution $\tau$ on $X$ extends naturally to an involution $\tau^{+}$ on $X^{+}$, which satisfies $\tau^{+}(+)=+$. Thus $KR^{0}(X,\tau)$ is defined to be the kernel of
$\ker\left[KR^{0}(X^{+},\tau^{+})\xrightarrow{\iota_{*}}KR^{0}(+,\mathrm{id})\right]\;.$
@@
As far as we know, the product structure on $A=\real(X,\tau)$ with $\tau$ non-trivial has not been investigated. If we identify $KO_{*}(\real(X,\tau))$ with Atiyah’s Real $K$-theory $KR^{*}(X,\tau)$, then there is a natural coproduct
@@
that induces a product structure on $KO_{*}(\real(X,\tau))\cong KR^{*}(X)$ exactly analogous to the product structures on $KO^{*}(X)$ and $K^{*}(X)$. This structure needs more attention than we can give it at this time.
@@
17. Clifford Algebras: Clifford and Lie algebra
The classic reference for Clifford algebras and their connection to $K$-theory is Atiyah, Bott, and Shapiro [9]. Husemoller [44] has a fine treatment. For connections with physics we rely on Baez [25], Stone et.al. [81] and G. Moore [64]. Notation for Clifford algebras varies a lot. When in doubt, please consult the chart below, which gives our notation.

### v3

@@ 14. The Basics: $KO$-theory for Spaces
For a compact space $X$, define
@@
    so, in particular, $TS^{n}$ plus a trivial bundle is isomorphic to a trivial bundle, whether $TS^{n}$ is trivial or not. Such a bundle is called stably trivial. So the tangent bundle of $S^{n}$ is stably trivial for all $n$. This implies $TS^{n}]=[\theta_{n}]$ in $KO^{0}(S^{n})$ for all $n$.
    It is a rather deep result that $TS^{n}$ is isomorphic to a trivial bundle if and only if n=1, 3,n=1,\,3, or $7$. (This is related to the existence of the complex numbers, the quaternions, and the Cayley numbers in dimensions 2, 4,2,\,4, and $8$ and the fact that there are no other such algebras in other dimensions.)
@@
$Vect^{\mathbb{R}}(X)\simeq\mathcal{P}_{\infty}(C^{\mathbb{R}}(X))$
between the abelian semigroup $Vect^{\mathbb{R}}(X)$ of isomorphism classes of vector bundles over $X$ and the abelian semigroup $\mathcal{P}_{\infty}(C^{\mathbb{R}}(X))$ of projections in matrix algebras over $C^{\mathbb{R}}(X)$. This induces a natural isomorphism of functors
@@
called Chern classes (cf. Milnor-Stasheff [63], Chapter 14). These may be defined using differential geometry or purely topologically. Here’s a quick sketch. Suppose $E\to X$ is a complex bundle of dimension $n$. We use the fact that the integral cohomology of the space $BU_{n}$ is a polynomial algebra over the integers with chosen generators $\{c_{1},\dots,c_{n}\}$ with $dim(c_{j})=2j$:
@@
    $ch:K^{*}(X)\otimes{\mathbb{Q}}\overset{\cong}{\longrightarrow}\bigoplus_{k}H^{k}(X;{\mathbb{Q}}).$
@@
Moving on, we come to the case of $KO^{*}(X)$, and for this we refer to Milnor-Stasheff [63] Chapter 15 and to Hatcher [40]. Suppose that $E\to X$ is a real vector bundle with complexification $E\otimes{\mathbb{C}}$. It turns out that
@@
    $Ph:KO^{*}(X)\longrightarrow\bigoplus_{k}H^{k}(X;{\mathbb{Q}}).$
@@
(Grady and Sati [38], §2.1.) It’s interesting to look at the simplest case, when $X$ is a point $pt$. Then $KO^{-k}(pt)=KO_{k}({\mathbb{R}})$ and so $KO^{*}(pt)\otimes{\mathbb{Q}}={\mathbb{Q}}$ when =0,4*=0,4 and is 0 otherwise. The generator of $KO^{0}(pt)$ is the identity of the ring, and it is represented by a one-dimensional trivial bundle over the point. It goes to the generator of $H^{0}(pt;{\mathbb{Z}})\cong{\mathbb{Z}}\subset H^{0}(pt;{\mathbb{Q}})$ since $Ph$ is a ring map. The group $KO^{-4}(pt)$ is more interesting. We have a natural isomorphism
@@
15. Relating $KO_{*}$-theory for $C^{*}$-Algebras and $KO^{*}$-theory for Spaces
@@
Suppose that $X$ is a compact space. Then the map $\Theta$ is an isomorphism.
@@
defines an injective map of vector bundles from $E$ to $X\times{\mathbb{R}}^{nk}$. Using the standard inner product on ${\mathbb{R}}^{nk}$, for each $x\in X$, let $p_{x}$ be the orthogonal projection from ${\mathbb{R}}^{nk}$ to the range of $f$ over $x$. Then $p_{x}$ is a continuous function of $x$ and is a projection in the algebraic sense that $p^{2}=p$ and $p^{*}=p$. Thus $p]\in\mathcal{P}_{\infty}(A)$ and by construction we have $\Theta([p])\cong[E]$. ∎
@@
$KO^{0}(X)=\ker\left[KO^{0}(X^{+})\xrightarrow{\iota^{*}}KO^{0}(+)\right]\;$ (15.1)
@@
$KO_{0}(A)=\ker\left[KO_{0}(A^{+})\xrightarrow{\pi_{*}}KO_{0}({\mathbb{R}})\right]\;$ (15.2)
@@
specifically evaluation at the point ++. This evaluation map is dual to the inclusion map $\iota$. This shows that the maps shown in Equations 15.1 and 15.2 are identical and hence there is an isomorphism $KO^{0}(X)\cong KO_{0}(A)$.
@@
Note that in Part (5) above, $KR^{0}(~)$” refers to Atiyah’s “Real $K$-theory” which is described in the next section.
@@ 16. Atiyah’s $KR$- theory
This is a contravariant functor on the category of topological spaces with involution. We assert that $KR^{*}$-theory for spaces with involution can be expressed in terms of $KO_{*}$- theory for an associated real $C^{*}$-algebra. Here is the precise statement.
@@
One should compare Karoubi [55] Theorem 4.1 for a related result. Before beginning the proof, we establish a basic property of complex involutive bundles over spaces with involution.
Note that the trivial bundle over $(X,\tau)$ of dimension $n$ is given by $(X\times{\mathbb{C}}^{n},\tau\times\psi_{n})$ where $\psi_{n}$ is pointwise conjugation on ${\mathbb{C}}^{n}$.
@@
Let $(X,\tau)$ be a compact space with involution and let $(E,\sigma)$ be a complex vector bundle over $X$ with a conjugate linear involution satisfying $\sigma(E_{x})=E_{\tau(x)}$. Then $(E,\sigma)$ is isomorphic (as a complex vector bundle with involution) to a sub-bundle $(E^{\prime},\sigma)$ of the trivial bundle $(X\times{\mathbb{C}}^{n},\tau\times\psi_{n})$ for some $n$. Further, there is a subbundle $(F,\rho)$ of the trivial bundle such that
$(E^{\prime},\sigma^{\prime})\oplus(F,\rho)\cong(X\times{\mathbb{C}}^{n},\tau\times\psi_{n})$
@@
First we work locally with the claim that for all $x_{0}\in X$ there exists a neighborhood $\mathcal{U}$ of $x_{0}$ in $X$ such that $\tau(\mathcal{U})=\mathcal{U}$ and $(E_{\mathcal{U}},\sigma)$ is isomorphic to $(\mathcal{U}\times{\mathbb{C}}^{n},\tau\times\psi_{n})$
Let $x_{0}\in X$ be given and let $\mathcal{V}$ be a neighborhood of $x_{0}$ on which $E$ is trivial. If $\tau(x_{0})\neq x_{0}$ we can arrange without loss of generality that $\mathcal{V}\cap\tau(\mathcal{V})=\emptyset$. Let $\phi\colon E_{\mathcal{V}}\rightarrow\mathcal{V}\times{\mathbb{C}}^{n}$ be an isomorphism; we use the notation $\phi_{x}\colon E_{x}\rightarrow{\mathbb{C}}^{n}$ for the map on the fibres. Let $\mathcal{U}=\mathcal{V}\cup\tau(\mathcal{V})$ and extend $\phi$ to
$\phi\colon E_{\mathcal{U}}\rightarrow\mathcal{U}\times{\mathbb{C}}^{n}$
by defining
$\phi(x,v)=(x,\psi(\phi_{\tau(x)}(\sigma_{x}(v))))\;\quad\text{for $x\in\tau(\mathcal{V})$.}$
Now $\phi\colon E_{\mathcal{U}}\rightarrow\mathcal{U}\times{\mathbb{C}}^{n}$ is defined and satisfies $(\tau\times\psi_{n})\circ\phi=\phi\circ\sigma$.
@@
Since $\widetilde{\sigma}_{x}$ is conjugate-linear and depends continuously on $x$, then $\psi_{n}\circ\widetilde{\sigma}$ is linear and can be represented an $n\times n$ matrix $A_{x}$ depending continuously on $x$. Thus we can write
$\widetilde{\sigma}_{x}=\psi_{n}\circ A_{x}\;,$
conflating the matrix $A_{x}$ with the linear transformation on ${\mathbb{C}}^{n}$ that it represents. The condition that $\widetilde{\sigma}$ is an involution implies that ${A_{x}}\overline{A_{\tau(x)}}=I_{n}$ for all $x\in\mathcal{U}$. Indeed, $(\widetilde{\sigma})^{2}=\mathrm{id}$ can be rewritten as
$\psi_{n}A_{\tau(x)}\psi_{n}A_{x}v=v\qquad\text{for}\qquad v\in{\mathbb{C}}^{n}$
@@
The isomorphism $\theta$ that we are looking for is from $\mathcal{U}\times{\mathbb{C}}^{n}$ to itself, so it can be represented by a continuous matrix-valued function $B_{x}$ for $x\in\mathcal{U}$. We need $\theta$ to carry the involution $\widetilde{\sigma}$ to the involution $\psi_{n}$ on ${\mathbb{C}}^{n}$. In other words, we need the diagram
@@
$(B_{\tau(x)}\circ\psi_{n}\circ A_{x})(v)=(\psi_{n}\circ B_{x})(v).$
@@
There is a continuous square root defined on square matrices in a neighborhood of the identity. We replace $\mathcal{U}$ with a smaller neighborhood to insure that square roots are defined on $A_{x}$ for $x\in\mathcal{U}$, and let $B_{x}=(A_{x})^{1/2}$. Then using the formula $A_{\tau(x)}=(\overline{A_{x}})^{-1}$ we have $B_{\tau(x)}=(\overline{A_{x}})^{-1/2}$ so
@@
$(E,\sigma)\longrightarrow(X\times{\mathbb{C}}^{m},\tau\times\psi_{n})$
@@
Finally, take $(E,\sigma)=(E,\tau\times\psi^{n})$ to be a subbundle of the trivial bundle. Then take $(F,\sigma)=(F,\tau\times\psi^{n})$ to be the point-wise orthogonal complement to $E$. Check easily that since vector conjugation on ${\mathbb{C}}^{n}$ respects the inner product, we have $(\tau\times\psi^{n})(F)=F$. Then we clearly have
$(E,\tau\times\psi^{n})\oplus(F,\tau\times\psi^{n})=(X\times{\mathbb{C}}^{n},\tau\times\psi^{n})$
completing the proof. ∎
The rest of this section is devoted to a new proof of Theorem 16.1, using Lemma 16.2. Let $(X,\tau)$ be a compact space with involution and let $A=\real(X,\tau)$.
We shall show that there is a natural isomorphism of abelian semigroups
@@
Proof of Theorem 16.1.
First let $p\in M_{n}(A)$ be a projection. Then $p\in C(X,M_{n}({\mathbb{C}}))$ and $p$ satisfies
$p(\tau(x))=\overline{p(x)}\;$
where the conjugation is entry-wise conjugation on matrices. Let
@@
with the associated projection map $\pi\colon E\rightarrow X$ by $\pi(x,v)=x$. Now $E$ is a complex vector bundle and has a conjugate-linear involution $\sigma=\tau\times\psi_{n}$ given by $(x,v)\mapsto(\tau(x),\overline{v})$. This involution $\sigma=\tau\times\psi_{n}$ certainly commutes with the respective involutions on $X$ as required. To show that $(\tau(x),\overline{v})\in E$ we check that
@@
We must show that $\Theta$ is well-defined on $\mathcal{P}_{\infty}$. Suppose that $p$ and $q$ are projections in $M_{n}(A)$ and that $u\in M_{n}(A)$ is a unitary that satisfies $upu^{*}=q$. Let
@@
Furthermore, it is straightforward to see that $p$ is a projection in $M_{n}(A)$, and if $p^{\prime}=p\oplus 0\in M_{n+m}(A)$, then $E_{p}\cong E_{p^{\prime}}$. More generally, if $p\in M_{m}(A)$ and $q\in M_{n}(A)$ are projections, then $p\oplus q$ is a projection in $M_{m+n}(A)$ and there is an isomorphism
$E_{p\oplus q}\cong E_{p}\oplus E_{q}\;.$
Thus the construction $p\mapsto(E,\sigma)$ is well-defined and gives a homomorphism of abelian semigroups
@@
Next, we claim that $\Theta$ is injective. Assume that $p,q$ are projections in $M_{n}(A)$ that satisfy $E_{p}\cong E_{q}$ (again as complex vector bundles with involution over $(X,\tau)$). Then there exists an element $a\in C(X,M_{n}({\mathbb{C}}))$ that implements this isomorphism (as in the first paragraph of the proof of Lemma 15.4). So $a_{x}$ is an isomorphism from $p(x){\mathbb{C}}^{n}$ to $q(x){\mathbb{C}}^{n}$ for each $x\in X$. Furthermore, since the isomorphism commutes with the involutions on $E_{p}$ and $E_{q}$, we have $a_{\tau(x)}(\overline{v})=\overline{a_{x}(v)}$ for all $x\in X$ and all $v\in\pi^{-1}(x)$. This implies that $a_{\tau(x)}(\overline{v})=\overline{a_{x}}\,\overline{v}$ and hence $a_{\tau(x)}=\overline{a_{x}}$ for all $x\in X$. Therefore $a\in M_{n}(A)$.
As in the proof of Lemma 15.4, we let $s=f(a^{*}a)a^{*}$ where
@@
noting in particular that because $a\in M_{n}(A)$, we also have $s\in M_{n}(A)$. As before, $s$ is a partial isometry that satisfies $ss^{*}=p$ and $s^{*}s=q$. Since
@@
we have again that $\mathrm{diag}(p,0)$ is unitarily equivalent to $\mathrm{diag}(q,0)$ in $M_{2n}(A)$, so $[p]=[q]$ in $\rm{Proj}(A)$. Thus $\Theta$ is injective.
Next, we show that $\Theta$ is surjective. Let $(E,\sigma)$ be any complex vector bundle with involution over $(X,\tau)$. Using Lemma 16.2, identify $E$ with a sub-bundle of $X\times{\mathbb{C}}^{n}$ in such a way that the involution $\sigma$ on $E$ corresponds to conjugation on ${\mathbb{C}}^{n}$. In other words, $\sigma$ can be expressed by the formula
@@
where $\overline{v}=\psi_{n}(v)$ is component-wise conjugation on ${\mathbb{C}}^{n}$. Define $p=p(x)\in C(X,M_{n}({\mathbb{C}}))$ by letting $p(x)$ be the orthogonal projection of ${\mathbb{C}}^{n}$ to $E_{x}$. Then $p$ is continuous in $x$, and is a projection: $p^{2}=p^{*}=p$.
Now, let $v\in E_{x}$. This implies that $\overline{v}\in E_{\tau(x)}$. Therefore we have $p(x)\cdot v=v$ and $p({\tau(x)})\cdot\overline{v}=\overline{v}$. This implies also that
$\overline{p({\tau(x)})}\cdot v=\overline{p{(\tau(x))}\cdot\overline{v}}=\overline{\overline{v}}=v\;.$
As $\overline{p(\tau(x))}$ and p(x){p(x)} are both orthogonal projections with the same rank which agree on $E_{x}$, they must be equal, implying that $\overline{p(\tau(x))}={p(x)}$ and hence $p(\tau(x))=\overline{p(x)}\;$. From this it follows that $p$ is an element in $M_{n}(A)$. By construction we have $(E,\sigma)\cong E_{p}$, which shows that $\Theta$ is surjective.
@@
The proof in the locally compact case follows the same strategy as that in the proof of Theorem 15.2. Note that if $X$ is locally compact then any involution $\tau$ on $X$ extends naturally to an involution $\tau^{+}$ on $X^{+}$, which satisfies $\tau^{+}(+)=+$. Thus $KR^{0}(X,\tau)$ is defined to be
$KR^{0}(X,\tau)=\ker\left[KR^{0}(X^{+},\tau^{+})\xrightarrow{\iota_{*}}KR^{0}(+,\mathrm{id})\right]\;.$
@@
Since $(\real(X,\tau))$ is commutative, there is a product structure on $KO_{*}(\real(X,\tau))$. As far as we know this product structure has not been investigated, except in the case with $\tau=\mathrm{id}$. If we identify $KO_{*}(\real(X,\tau))$ with Atiyah’s Real $K$-theory $KR^{*}(X)$, then there is a natural coproduct
@@
that induces a product structure on $KO_{*}(\real(X,\tau))\cong KR^{*}(X)$ exactly analogous to the product structures on $KO^{*}(X)$ and $K^{*}(X)$. This needs more attention than we can give it at this time.
@@
17. Clifford Algebras and Their Associated Lie Groups
The classic reference for Clifford algebras and their connection to $K$-theory is Atiyah, Bott, and Shapiro [9]. We refer to this paper and these authors by ABS. Husemoller [44] has a fine treatment. For connections with physics we rely on Baez [25], Stone et.al. [81] and G. Moore [64]. Notation for Clifford algebras varies a lot. When in doubt, please consult the chart below, which gives our notation.